OpenStudy (anonymous):

f(x) = -2sin(x-pi/2) what are two x intercepts

OpenStudy (anonymous):

sine is zero if the input is zero, so you can start by solving \[x-\frac{\pi}{2}=0\] and solve for \(x\) in one step

OpenStudy (anonymous):

sine is also zero if \(x=\pi\) so you can then solve \[x-\frac{\pi}{2}=\pi\] again in one step

OpenStudy (anonymous):

0 and pi/2

OpenStudy (anonymous):

yes